Write and solve a system of equations for a total of $15,000 invested in two accounts, with simple interest rates 1.85% and 2.95% respectively, where the total interest earned was $425.
Answer by mananth(16949)

Part I 1.85% per annum ------------- Amount invested =x
Part II 2.95% per annum ------------ Amount invested = y

Interest----- 425

Part I 1.85% per annum ---x
Part II 2.95% per annum ---y
Total investment
x + 1 y= 15000 -------------1
Interest on both investments
1.85% x + 2.95% y= 425
Multiply by 100
1.85 x + 2.95 y= 42500.00 --------2
Multiply (1) by -1.85
we get
-1.85 x -1.85 y= -27750.00
Add this to (2)
0 x 1.10 y= 14750
divide by 1.1
y = 13409.09
Part I 1.85% $ 1590.91
Part II 2.95% $ 13409.09

CHECK
1590.91 --------- 1.85% ------- 29.43
13409.09 ------------- 2.95% ------- 395.57
Total -------------------- 425.00

m.ananth@hotmail.ca